OpenStudy (anonymous):

Kenny goes shopping for school clothes during a tax−free weekend, and he saves eight percent sales tax. Use a percent proportion to determine how much he would have to spend to save ten dollars. Which of the following is true?

OpenStudy (anonymous):

The solution is eighty dollars, and is a denominator in the proportion. The solution is eighty dollars, and is a numerator in the proportion. The solution is one hundred twenty-five dollars, and is a denominator in the proportion. The solution is one hundred twenty-five dollars, and is a numerator in the proportion.

Directrix (directrix):

@nyprincess2001 I am going to post my thoughts on this but percent proportion is something about which I am not sure. Here we go.

Directrix (directrix):

Let P is the total tax-free purchases bought by Kenny. Kenny wants to spend $P with no tax and then be able to say he saved exactly $10 of tax money. So, 8% of $P = $10 8% = 8/100 = .08 .08 P = 10 P = $125 BUT what about the percent proportion options? @nyprincess2001

Directrix (directrix):

Oh, I did not use the proportion method.

Directrix (directrix):

This --> 8% of $P = $10 could be written as: (8*P)/100 = 10/1 which is a proportion. @nyprincess2001 Let me know what you think.